OpenStudy (anonymous):

The ball launcher in a pinball machine has a spring that has a force constant of 1.22 N/cm. The surface on which the ball moves is in- clined 15.8 with respect to the horizontal. If the spring is initially compressed 3.68 cm, find the launching speed of a 0.172 kg ball when the plunger is released. The acceleration due to gravity is 9.8 m/s2 . Friction and the mass of the plunger are negligible. Answer in units of m/s

OpenStudy (agent0smith):

Work done = integral of spring force minus component of gravity down incline \[\Large \int\limits(-kx -0.172*9.8 \sin 15.8)dx\] hmmm after writing that, maybe you don't need an integral. Just find the net work done (work is force times distance: Work done by spring = 1/2*kx^2 where x=0.0368 m (metres not cm) work done by gravity = 0.172*9.8*sin15.8*x where x=0.0368m convert k=1.22 N/cm to 122N/m and calculate net work (spring work minus gravity) Net work done = 1/2*122(0.0368)^2 - 0.172*9.8*sin15.8*0.0368 =0.08831 Joules

OpenStudy (agent0smith):

work done = change in KE = 1/2*mv^2 where m = 0.172 gives v = 1.01 m/s

OpenStudy (anonymous):

I understand how you got that answer but it is incorrect for this problem

OpenStudy (agent0smith):

Hmm, idk then :(

OpenStudy (anonymous):

It's ok :) it's a sucky problem lol

OpenStudy (agent0smith):

I'll try to figure it out, maybe i have a number wrong... i'm pretty sure the method is right.

OpenStudy (anonymous):

Sounds good I've tried different wAys to get it but it hasn't worked

OpenStudy (agent0smith):

Do you know what the answer is?

OpenStudy (agent0smith):

this problem is solved EXACTLY the same way i solved your problem: http://www.csupomona.edu/~skboddeker/131/131hw/ch7h.htm scroll down to: Ch 7 #63

OpenStudy (anonymous):

I don't know what the answer is :/ But ya that's what I'm saying I've tried

OpenStudy (agent0smith):

So at this point i wonder if the "correct" answer to the problem is wrong... if you and i have tried multiple times, and my method is exactly the same as the one above... I'm checking all my numbers just to be sure but it looks all correct.

OpenStudy (agent0smith):

Ask your lecturer or whoever...

OpenStudy (anonymous):

Ha ha imma have to thanks for your help though :)

OpenStudy (agent0smith):

No prob. Let me know what you find out... cos i can't find any errors, and i've looked up a few similar questions, all solved the same.

OpenStudy (agent0smith):

@RavenLynette i found the error haha... i used wolfram alpha to find the work done, and it used radians not degrees!

OpenStudy (agent0smith):

1/2*122(0.0368)^2 - 0.172(9.8)(0.0368*sin(15.8)) Net work done = 1/2*122(0.0368)^2 - 0.172*9.8*sin15.8*0.0368 =0.0657191 Joules (using degrees this time not radians...) work done = change in KE = 1/2*mv^2 = 0.0657191 v = 0.874 m/s

OpenStudy (anonymous):

Haha omg thank you :3

OpenStudy (agent0smith):

you're welcome! :D